OpenStudy (anonymous):

23. Name three faces of the cube above that intersect at point G. (1 point) a. DHGC, CBFG, GFEH b. ABFE, ADHE, ABCD c. DCBA, GEBF, GFEH d. GDAF, CBFG, GFEH 26. Find the surface area of the cylinder to the nearest tenth of a square unit. Use pi. = 3.14. a. 206 cm2 b. 100.5 cm2 c. 623 cm2 d. 412 cm2 mc0262.jpg nar0021.jpg

OpenStudy (anonymous):

I'll give you one face for the cube, and you should be able to find two more. One is ABCD.

OpenStudy (anonymous):

As for the surface area of the cylinder.. do you know the formula to find the surface area of a cylinder?

OpenStudy (anonymous):

no..

OpenStudy (anonymous):

which unit and lesson

OpenStudy (anonymous):

unit 2 lesson 16. 7th grade

OpenStudy (anonymous):

Here's the formula for the surface area of a cylinder: \[2\pi r ^{2} + 2\pi r h\] Where r = radius h = height \[\pi = p i\]and pi = 3.14

OpenStudy (anonymous):

Do you know how to plug your values in?

OpenStudy (anonymous):

ohh im in 10th grade lol

OpenStudy (anonymous):

No....

OpenStudy (anonymous):

i tried but i got a WAY different answer then the answers provided..

OpenStudy (anonymous):

What values did you plug into where? (what did you use for the radius and what did you use for the height?)

OpenStudy (anonymous):

wait i think i got it is it 412???

OpenStudy (anonymous):

Well, here's how it should look before solving:\[(2)(3.14)(4^{2}) + (2)(3.14)(4)(12.4)\]

OpenStudy (anonymous):

If you simplify that by squaring the radius, then it's: (2)(3.14)(16) + (2)(3.14)(4)(12.4) How would you go bout solving this?

OpenStudy (anonymous):



OpenStudy (anonymous):

Yes! Very good!

OpenStudy (anonymous):

Thx!!